(a) Use Fermat’s little theorem to compute: 4101 mod 5, 4101 mod 7, 4101 mod 11.


(b) Use your results from part (a) and the Chinese Remainder Theorem to compute


4


101 mod 385. (note that 385 = 5 × 7 × 11).

1
Expert's answer
2021-11-03T07:57:32-0400

a) "4^{101} mod 5"

"4^2mod5=1"

"4^{101}mod5=4^{100+1}mod5=(4*4^{2*50})mod5=(4*(4^{2})^{50})mod5=4*1^{50}=4"

b) "4^{101} mod 7"

"4^3mod7=1"

"4^{101}mod7=4^{100+1}mod7=(4^2*4^{3*33})mod7=(4^2*(4^{3})^{33})mod7=(4^2*1^{33})mod7=2"

c) "4^{101} mod 11"

"4^5 mod 11=1"

"4^{101}mod11=4^{100+1}mod11=(4*4^{5*20})mod11=(4*(4^{5})^{20})mod11= 4*1^{20}=4"


We have that "M{\scriptscriptstyle 1}={\frac {385} 5}=77" , and the inverse of 77 modulo 5 is y1 = 3. As well, "M{\scriptscriptstyle 1}={\frac {385} 7}=55" , and the inverse of 55 modulo 7 is y2 = 6. Finally, "M{\scriptscriptstyle 1}={\frac {385} {11}}=35" , and the inverse of 35 modulo 11 is y3 = 6. Thus, by the Chinese Remainder Theorem, the solution has the form:

"4^{101}mod385= (4*77*3(mod385) +2*55*6(mod385) +4*35*6(mod385))" mod385 =

= 114